# Robust Photo-Realistic Hand Gesture Generation: from Single View to Multiple View

### v1

@@ Abstract
High-fidelity hand gesture generation represents a significant challenge in human-centric generation tasks. Existing methods typically employ single-view 3D MANO mesh-rendered images prior to enhancing gesture generation quality. However, the complexity of hand movements and the inherent limitations of single-view rendering make it difficult to capture complete 3D hand information, particularly when fingers are occluded. The fundamental contradiction lies in the loss of 3D topological relationships through 2D projection and the incomplete spatial coverage inherent to single-view representations. Diverging from single-view prior approaches, we propose a multi-view prior framework, named Multi-Modal UNet-based Feature Encoder (MUFEN), to guide diffusion models in learning comprehensive 3D hand information. Specifically, we extend conventional front-view rendering to include rear, left, right, top, and bottom perspectives, selecting the most information-rich view combination as training priors to address occlusion completion. This multi-view prior with a dedicated dual stream encoder significantly improves the model’s understanding of complete hand features. Furthermore, we design a bounding box feature fusion module, which can fuse the gesture localization features and gesture multi-modal features to enhance the location-awareness of the MUFEN features to the gesture-related features. Experiments demonstrate that our method achieves state-of-the-art performance in both quantitative metrics and qualitative evaluations.
@@
Figure 1: Photo-Realistic Hand Gesture Generation via various methods. Our method (fifth row) improves realism and accuracy by fusing multi-view and multi-modal features. Results are closer to ground truth (bottom row).
1 Introduction
@@
Human-centric image and video generation has been a long-standing focus in the field of computer vision [27, 14]. Despite remarkable advancements in diffusion models that have enabled more realistic human generation [23] and improved pose control accuracy [26, 10], achieving natural and precise hand gesture generation remains a critical challenge, with persistent issues of unnatural occurrences such as finger multiplicity or deficiency, as shown in Figure 1. This challenge primarily arises from two inherent characteristics: Firstly, the hand is highly flexible, capable of diverse gestures - although occupying minimal spatial proportion relative to the full body, the hands possess extraordinary degrees of freedom and movement complexity. Secondly, severe occlusion phenomena, particularly inter-finger occlusions, create substantial obstacles for models to learn comprehensive hand representations, potentially leading to incomplete hand modelling through partial observation.
@@
The challenge of diverse hand gestures can be addressed through the enhancement of dataset content, model structure, and training algorithm. Upper-body-focused datasets are carefully designed to emphasize hand gesture representations [8, 11, 3]. Notably, HaGRID [8] – originally designed for gesture recognition – has emerged as the most popular dataset for hand generation tasks. Beyond specialized gesture dataset construction, numerous methods have developed dedicated architectures or algorithms to enhance gesture generation quality. These approaches are broadly categorized into single-stage generation, and multi-stage inpainting methodologies. For single-stage generation, researchers primarily design dedicated encoders to extract rich features from 3D meshes or depth image priors, thereby guiding generation models to synthesize high-fidelity hand gestures [16]. Multi-stage approaches typically employ specialized generative models to refine hand regions after initial full-image generation by general-purpose models, implementing targeted corrections through cascaded networks [13].
@@
Although existing methods have made progress in improving the quality of gesture generation, they typically rely on mesh renderings from a single viewpoint as prior [13, 23]. This limitation prevents the model from acquiring sufficient 3D information about the hand gestures, especially in cases of self-occlusion, where critical structural information is often missing. As illustrated in Figure 2 with the thumb obscured, complex gestures often involve severe occlusions, particularly in the fingers, where occluded parts result in ambiguities, making it difficult for the model to infer accurate finger configurations. Such inaccurate modelling severely hinders the learning of complete hand representations, leading to significant distortions during gesture generation, such as missing or multiple fingers. To address this issue, we enhance the 3D representation of gestures through multi-view feature fusion, providing the model with more comprehensive hand modelling information for more accurate gesture synthesis.
@@
To address the occlusion issues in gestures and enable the model to learn comprehensive hand modelling information, we rendered the meshes from frontal, rear, left, right, top, and bottom perspectives to supplement the occlusion information in a single view. We then designed an encoder with shared parameters for the rendering meshes from different perspectives to extract visual features from these views. Additionally, to enhance the spatial information of multimodal hand features extracted by MUFEN, a bounding box feature fusion module is proposed for enhancing multimodal gesture features with gesture localization features. This allows the MUFEN features to be fused directly with the features of the diffusion model, without the need for separate adjustments to the fusion region like zero padding strategies [4] that have appeared in the literature.
@@
  • We utilize multi-view (including frontal, rear, left, right, top, and bottom perspectives) gesture 3D mesh priors to supplement occluded hand information in single-view scenarios.
@@
  • We design a multimodal encoder based on the gesture region localization information to extract and fuse the multimodal features of the gesture region, to enhance the model’s ability to learn comprehensive hand modelling information.
@@
  • We leverage gesture localization features from a bounding box to enhance MUFEN features, enabling them to be fused directly with the diffusion model features without the need for fusion region adjustment.
@@
  • Extensive simulations demonstrate that the proposed method achieves state-of-the-art (SOTA) performance in both quantitative and qualitative analyses.
2 Related Work
2.1 Controllable Human-Centric Generation
@@
Controllable human-centric image or video generation is a popular task. Recently, ControlNet [26] and T2I-Adapter [15] enable conditional human pose generation by incorporating control signals such as skeletons, sketches, and depth maps. And then, ControlNet++ [10] improved control precision under multiple conditions by introducing a pixel-level cycle consistency loss, enhancing realism. Follow Your Pose [14] fine-tuned a diffusion model with human-centric data in the first stage and trained temporal attention layers in the second stage to achieve temporal consistency, enabling pose-controllable human video generation. Building on this, Follow Your Pose v2 [25] incorporated multiple control signals, such as depth maps and optical flow, to improve the accuracy of pose control and temporal consistency.
2.2 Photo-Realistic Hand Gesture Generation
@@
Recent approaches aiming at improving the fidelity of gesture generation can generally be grouped into two categories: single-stage methods [16, 18, 5] and multi-step pipelines [13, 20, 23, 27].
@@
Single-stage Methods: These methods produce the entire image in one forward pass while simultaneously directing the model’s focus toward learning the detailed features of the hand area. For instance, HandDiffuser [16] strengthens the diffusion model’s ability to capture the structural patterns of hand gestures by integrating 3D hand mesh data through a dedicated encoding module. Also leveraging the rich information from 3D mesh, a novel text-guided framework named AttentionHand [18] uses 3D hand mesh prior with text and bounding box for controllable hand image generation. Overall, achieving realistic hand generation within a one-step framework is still a complex task due to data-intensive and the need to simultaneously maintain both holistic image coherence and intricate hand geometry.
@@
Multi-stage Methods: These approaches divide the gesture generation task into several stages, where each stage applies a specific generative model to handle particular regions of the image. The outputs from all stages are then combined into a final result. A key step in this pipeline is localized inpainting: based on the coarse image from the initial stage, a hand-focused inpainting model is used to enhance the quality of the hand area. Compared with single-stage frameworks, multi-stage strategies can improve the realism and accuracy of hand synthesis, although they usually require more complex inference procedures and additional model components. HandRefiner [13] adapts ControlNet [26] by fine-tuning it with hand depth maps, allowing the inpainting diffusion model to repair the distorted hand area. However, it still struggles to maintain consistency between the refined hand area and the surrounding content generated in the first step. Leveraging hand and face masks from DINO v2 [17] for segmentation, RealisHuman [23] introduces a diffusion model controlled by multiple prior conditions to generate realistic facial and hand areas, followed by a second step to repaint the transitions between background and modified human parts. Nevertheless, the prior used comes from a 2D projection of a single viewpoint and still does not build complete structural information. Although previous single-stage and multi-stage approaches leverage multi-modal priors to enhance the quality of hand gesture generation, none of them incorporate complementary information from multiple viewpoints. This is particularly important for hand gestures, which often involve complex spatial structures and severe occlusions. In contrast to these methods, our proposed MUFEN extracts and fuses multi-view features along with a multi-modal feature of hand gestures. By providing a more comprehensive representation of hand gestures, our method guides the diffusion model to generate more realistic hand gestures within a single-stage inference framework, removing the need for additional stages or external inpainting processes.
@@
Figure 2: Multi-view hand MANO mesh rendering pipeline.
3 MUFEN
3.1 Preliminaries
3.1.1 Latent Diffusion Model with ControlNet Guidance
@@
In this paper, we utilize a latent diffusion framework as our core generative backbone due to its strong generation performance and improved efficiency. Latent Diffusion Models (LDMs) [21] operate in a compressed latent space rather than directly in pixel space, significantly reducing computational requirements when handling high-dimensional data. During the forward diffusion stage, Gaussian noise is incrementally injected into the latent variable $\mathbf{z}_{0}$, producing a noisy version $\mathbf{z}_{t}$ according to the formula:
$\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
@@
$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon},t}\left[\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t)\|_{2}^{2}\right],$ (2)
@@
To further enhance the controllability of the generation process, especially for structure-aware image synthesis, ControlNet [26] is incorporated as an auxiliary module. ControlNet extends diffusion models by introducing an additional conditioning pathway that guides the generation according to specific structural cues, such as key points, depth, or semantic maps. This mechanism allows the model to follow external constraints while preserving the quality of the generated content. It does so by integrating control embeddings $c$ into the denoising network at every step of the diffusion. The objective function of ControlNet adapts the LDM loss into a conditional form:
$\mathcal{L}_{\text{denoise}}=\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon},t,c}\left[\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t,c)\|_{2}^{2}\right],$ (3)
@@
3.1.2 MANO Hand Representation
To model the 3D structure of hand gestures, we adopt the MANO model [22], which is a parametric hand model similar in design to the SMPL body model [12]. The MANO framework produces a three-dimensional hand mesh composed of 778 vertices, each described by a triplet of spatial coordinates $(x,y,z)$. Moreover, it provides a set of 21 semantic key points representing essential anatomical joints of the hand, including the wrist and four articulations per finger. For multi-view mesh rendering, we utilized the recent mesh annotator HaMeR  [19] for accurate recovery of MANO mesh parameters. This procedure allowed us to extract both the MANO pose and shape parameters for each hand gesture, along with the corresponding hand bounding boxes.
3.2 Multi-View MANO Mesh Rendering
Figure 2 illustrates our pipeline for rendering Multi-View MANO Meshes for each gesture. Specifically, for each input image, we first employ HaMeR [19] to extract the meshes of all hand gestures present in the image and render front-view mesh images of these gestures. Subsequently, we adjust the mesh vertex coordinates and camera positions through coordinate transformations to achieve rendering from other viewpoints. We begin by detailing the coordinate transformation from the front view to the rear view, which involves a 180-degree rotation around the Y-axis, effectively reversing the direction of observation. Specifically, the transformation from front view to rear view can be formalized as:

### v2

@@ Abstract.
High-fidelity hand gesture generation represents a significant challenge in human-centric generation tasks. Existing methods typically employ a single-view mesh-rendered image prior to enhancing gesture generation quality. However, the spatial complexity of hand gestures and the inherent limitations of single-view rendering make it difficult to capture complete gesture information, particularly when fingers are occluded. The fundamental contradiction lies in the loss of 3D topological relationships through 2D projection and the incomplete spatial coverage inherent to single-view representations. Diverging from single-view prior approaches, we propose a multi-view prior framework, named Multi-Modal UNet-based Feature Encoder (MUFEN), to guide diffusion models in learning comprehensive 3D hand information. Specifically, we extend conventional front-view rendering to include rear, left, right, top, and bottom perspectives, selecting the most information-rich view combination as training priors to address occlusion. This multi-view prior with a dedicated dual stream encoder significantly improves the model’s understanding of complete hand features. Furthermore, we design a bounding box feature fusion module, which can fuse the gesture localization features and multi-modal features to enhance the location-awareness of the MUFEN features to the gesture-related features. Experiments demonstrate that our method achieves state-of-the-art performance in both quantitative metrics and qualitative evaluations. The source code is available at https://github.com/fuqifan/MUFEN.
Multi-Modal Diffusion Model, Hand Gesture Generation, MANO mesh
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755828isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Reconstructionccs: Computing methodologies Visibility
@@
Figure 1. Photo-realistic hand gesture generation via various methods. Our method, named MUFEN, improves the realism and accuracy of gesture generation by fusing multi-view and multi-modal features of the gesture.
1. Introduction
@@
Human-centric image and video generation has been a long-standing focus in the field of computer vision (Zhu et al., 2024; Ma et al., 2024). Despite remarkable advancements in diffusion models that have enabled more realistic human generation (Wang et al., 2025) and improved pose control accuracy (Zhang et al., 2023; Li et al., 2024), achieving natural and precise hand gesture generation remains a critical challenge, with persistent issues of unnatural occurrences such as finger multiplicity or deficiency, as shown in Figure 1. This challenge primarily arises from two inherent characteristics: Firstly, the hand is highly flexible, capable of diverse gestures - although occupying minimal spatial proportion relative to the full body, the hands possess extraordinary degrees of freedom and movement complexity. Secondly, severe occlusion phenomena, particularly inter-finger occlusions, create substantial obstacles for models to learn comprehensive hand representations, potentially leading to incomplete hand modelling through partial observation.
@@
The challenge of diverse hand gestures can be addressed through the enhancement of dataset content, model structure, and training algorithm. Upper-body-focused datasets are carefully designed to emphasize hand gesture representations (Kapitanov et al., 2024; Lin et al., 2023; Duarte et al., 2021). Notably, HaGRID (Kapitanov et al., 2024) – originally designed for gesture recognition – has emerged as the most popular dataset for hand generation tasks. Beyond specialized gesture dataset construction, numerous methods have developed dedicated architectures or algorithms to enhance gesture generation quality. These approaches are broadly categorized into single-stage generation, and multi-stage inpainting methodologies. For single-stage generation, researchers primarily design dedicated encoders to extract rich features from 3D meshes or depth image priors, thereby guiding generation models to synthesize high-fidelity hand gestures (Narasimhaswamy et al., 2024). Multi-stage approaches typically employ specialized generative models to refine hand regions after initial full-image generation by general-purpose models, implementing targeted corrections through cascaded networks (Lu et al., 2024).
@@
Although existing methods have made progress in improving the quality of gesture generation, they typically rely on mesh renderings from a single viewpoint as prior (Lu et al., 2024; Wang et al., 2025). This limitation prevents the model from acquiring sufficient 3D information of the hand gestures, especially in cases of self-occlusion, where critical structural information is often missing. As illustrated in Figure 2 with the thumb obscured, complex gestures often involve severe occlusions, particularly in the fingers, where occluded parts result in ambiguities, making it difficult for the model to infer accurate finger configurations. Such inaccurate modelling severely hinders the learning of complete hand representations, leading to significant distortions during gesture generation, such as missing or multiple fingers.
@@
To address occlusion issues in gestures and enable the model to learn comprehensive hand modelling information, we rendered the meshes from front, rear, left, right, top, and bottom perspectives to supplement the occlusion information in a single view. We then designed a Multi-Modal UNet-based Feature Encoder (MUFEN) to extract multi-view multi-modal features of hand gesture. Specifically, MUFEN uses dual stream rendering encoder for the rendering meshes from different perspectives to extract multi-view features. Deep features and linguistic labeling features are extracted using a trainable deep encoder and a frozen CLIP, respectively. Additionally, to enhance the spatial information of multimodal hand features extracted by MUFEN, a bounding box feature fusion module is proposed for enhancing multimodal gesture features with gesture localization features. This allows the MUFEN features to be fused directly with the features of the diffusion model, without the need for separate adjustments to the fusion region like zero padding strategies (Fu et al., 2025) that have appeared in the literature.
@@
  • We utilize multi-view (including front, rear, left, right, top, and bottom perspectives) gesture rendering mesh priors to supplement occluded hand information in single-view scenarios.
@@
  • We design a multi-modal UNet-based feature encoder named MUFEN with dual stream rendering encoder to extract and fuse the multi-view multi-modal features of the gesture region, to enhance the model’s ability to learn comprehensive hand modelling information.
@@
  • We leverage gesture localization features from a bounding box to enhance MUFEN features by a trainable bounding box feature fusion module, enabling MUFEN features to be fused directly with the diffusion model features without the need for fusion region adjustment.
@@
  • Simulations demonstrate that the proposed MUFEN achieves state-of-the-art (SOTA) performance in both quantitative and qualitative analyses.
2. Related Work
2.1. Controllable Human-Centric Generation
@@
Controllable human-centric image or video generation is a popular task. Recently, ControlNet (Zhang et al., 2023) and T2I-Adapter (Mou et al., 2024) enable conditional human pose generation by incorporating control signals such as skeletons, sketches, and depth maps. And then, ControlNet++ (Li et al., 2024) improved control precision under multiple conditions by introducing a pixel-level cycle consistency loss, enhancing realism. Follow Your Pose (Ma et al., 2024) fine-tuned a diffusion model with human-centric data in the first stage and trained temporal attention layers in the second stage to achieve temporal consistency, enabling pose-controllable human video generation. Building on this, Follow Your Pose v2 (Xue et al., 2024) incorporated multiple control signals, such as depth maps and optical flow, to improve the accuracy of pose control and temporal consistency.
2.2. Photo-Realistic Hand Gesture Generation
@@
Recent approaches aiming at improving the fidelity of gesture generation can generally be grouped into two categories: single-stage methods (Narasimhaswamy et al., 2024; Park et al., 2024; Fu et al., 2024) and multi-step pipelines (Lu et al., 2024; Pelykh et al., 2024; Wang et al., 2025; Zhu et al., 2024).
@@
Single-stage Methods: These methods produce the entire image in one forward pass while simultaneously directing the model’s focus toward learning the detailed features of the hand area. For instance, HandDiffuser (Narasimhaswamy et al., 2024) strengthens the diffusion model’s ability to capture the structural patterns of hand gestures by integrating 3D hand mesh data through a dedicated encoding module. Also leveraging the rich information from 3D mesh, a novel text-guided framework named AttentionHand (Park et al., 2024) uses 3D hand mesh prior with text and bounding box for controllable hand image generation. Overall, achieving realistic hand generation within a one-step framework is still a complex task due to data-intensive and the need to simultaneously maintain both holistic image coherence and intricate hand geometry.
@@
Multi-stage Methods: These approaches divide the gesture generation task into several stages, where each stage applies a specific generative model to handle particular regions of the image. The outputs from all stages are then combined into a final result. A key step in this pipeline is localized inpainting: based on the coarse image from the initial stage, a hand-focused inpainting model is used to enhance the quality of the hand area. Compared with single-stage frameworks, multi-stage strategies can improve the realism and accuracy of hand synthesis, although they usually require more complex inference procedures and additional model components. HandRefiner (Lu et al., 2024) adapts ControlNet (Zhang et al., 2023) by fine-tuning it with hand depth maps, allowing the inpainting diffusion model to repair the distorted hand area. However, it still struggles to maintain consistency between the refined hand area and the surrounding content generated in the first step. Leveraging hand and face masks from DINO v2 (Oquab et al., 2024) for segmentation, RealisHuman (Wang et al., 2025) introduces a diffusion model controlled by multiple prior conditions to generate realistic facial and hand areas, followed by a second step to repaint the transitions between background and modified human parts. Nevertheless, the prior used comes from a 2D projection of a single viewpoint and still does not build complete structural information. Although previous single-stage and multi-stage approaches leverage multi-modal priors to enhance the quality of hand gesture generation, none of them incorporate complementary information from multiple viewpoints. This is particularly important for hand gestures, which often involve complex spatial structures and severe occlusions. In contrast to these methods, our proposed MUFEN extracts and fuses multi-view features along with a multi-modal feature of hand gestures. By providing a more comprehensive representation of hand gestures, our method guides the diffusion model to generate more realistic hand gestures within a single-stage inference framework, removing the need for additional stages or external inpainting processes.
@@
Figure 2. Multi-view hand MANO mesh rendering pipeline.
3. Multi-Modal UNet-based Feature Encoder
3.1. Preliminaries
3.1.1. Latent Diffusion Model with ControlNet Guidance
@@
In this paper, we utilize a latent diffusion framework as our core generative backbone due to its strong generation performance and improved efficiency. Latent Diffusion Models (LDMs) (Rombach et al., 2022) operate in a compressed latent space rather than directly in pixel space, significantly reducing computational requirements when handling high-dimensional data. During the forward diffusion stage, Gaussian noise is incrementally injected into the latent variable $\mathbf{z}_{0}$, producing a noisy version $\mathbf{z}_{t}$ according to the formula:
(1) $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$
@@
(2) $\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon},t}\left[\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t)\|_{2}^{2}\right],$
@@
To further enhance the controllability of the generation process, especially for structure-aware image synthesis, ControlNet (Zhang et al., 2023) is incorporated as an auxiliary module. ControlNet extends diffusion models by introducing an additional conditioning pathway that guides the generation according to specific structural cues, such as key points, depth, or semantic maps. This mechanism allows the model to follow external constraints while preserving the quality of the generated content. It does so by integrating control embeddings $c$ into the denoising network at every step of the diffusion. The objective function of ControlNet adapts the LDM loss into a conditional form:
(3) $\mathcal{L}_{\text{denoise}}=\mathbb{E}_{\mathbf{x},\boldsymbol{\epsilon},t,c}\left[\|\boldsymbol{\epsilon}-\epsilon_{\theta}(\mathbf{z}_{t},t,c)\|_{2}^{2}\right],$
@@
3.1.2. MANO Hand Representation
To model the 3D structure of hand gestures, we adopt the hand Model with Articulated and Non-rigid deformations (MANO) model (Romero et al., 2017), which is a parametric hand model similar in design to the SMPL body model (Loper et al., 2023). The MANO framework produces a three-dimensional hand mesh composed of 778 vertices, each described by a triplet of spatial coordinates $(x,y,z)$. Moreover, it provides a set of 21 semantic key points representing essential anatomical joints of the hand, including the wrist and four articulations per finger. For multi-view mesh rendering, we utilized the recent mesh annotator HaMeR  (Pavlakos et al., 2024) for accurate recovery of MANO mesh parameters. This procedure allowed us to extract both the MANO pose and shape parameters for each hand gesture, along with the corresponding hand bounding boxes.
3.2. Multi-View MANO Mesh Rendering
Figure 2 illustrates our pipeline for rendering Multi-View MANO Meshes for each gesture. Specifically, for each input image, we first employ HaMeR (Pavlakos et al., 2024) to extract the meshes of all hand gestures present in the image and render front-view mesh images of these gestures. Subsequently, we adjust the mesh vertex coordinates and camera positions through coordinate transformations to achieve rendering from other viewpoints. We begin by detailing the coordinate transformation from the front view to the rear view, which involves a 180-degree rotation around the Y-axis, effectively reversing the direction of observation. Specifically, the transformation from front view to rear view can be formalized as: